SBB09Ntbk-Nikhil Sharma

From OpenWetWare
Jump to: navigation, search

Nikhil Sharma 12:18, 23 February 2009 (PST)

Today, I digested my wobble product for Poly_bind basic part. First, I ran a small fragment zymo cleanup, followed by a digestion with EcoR1/BamH1, and then another zymo clean-up. For the TraT part, I was able to do an analytical gel to determine that my product was correct. I ran a normal zymo cleanup, but I ran out of time to run the digest. I froze my eluted DNA and will digest it on Wednesday.

Nikhil Sharma 12:31, 25 February 2009 (PST)

Today I took the cleaned-up wobble digest for the Poly_bind basic part and ligated it with its vector, pBca9495AK-Bca1144#5. After following the ligation protocol, I transformed the ligated product by heat shock and plated it onto an AK plate.

For the TraT part, I was able to digest and cleanup, but I did not have enough time to ligate it. I will take care of that in the next lab period.

Nikhil Sharma 13:31, 27 February 2009 (PST)

Today, I did a miniprep purification of the transformed Poly_bind basic part. I had enough colonies on my plate to suggest transformation had occurred. I wasn't able to do the mapping at the end of the purification, so I will do that on Monday.

I was able to transform the TraT part, and I will proceed to miniprep purify it on Monday.

Nikhil Sharma 14:05, 2 March 2009 (PST)

Today, I ran a miniprep purification of the TraT part. After miniprep purifying it, I mapped both the Poly_bind and TraT parts on a gel. After checking with Chris that what I had was correct, I placed my samples in the cloning plates to be taken for sequencing and filled out the sequencing log on the wiki. For now, I'm done with lab work.

Nikhil Sharma 10:50, 4 March 2009 (PST)

Today, I wasn't able to do anything related to the project, as my samples weren't sent in for sequencing. On Monday, Chris told me to fill out the sequencing GoogleDoc that is linked on bSpace, but I was sent to a different GoogleDoc when I clicked on the link. Apparently this is an error with bSpace? Regardless, because I didn't fill out the write document, my parts weren't sent in for sequencing, so today I will just work on the assignment due tomorrow.

Nikhil Sharma 11:20, 6 March 2009 (PST)

Today, I got my sequences back. I'll analyze them later to see if they are correct.

Nikhil Sharma 12:15, 9 March 2009 (PST)

There were some issues with the samples I sent in for sequencing. Roger and I apparently switched our samples earlier in the process (during the initial zymo cleanups) so he ended up sequencing my TraT part. The sequence I sent in for the polystyrene binding peptide was useless, as the part did not ligate correctly. Chris is going to send in my other sequence for the part.

Nikhil Sharma 13:13, 11 March 2009 (PST)

The other sequence I sent in for the {poly_bind} part came back correctly, so both of my parts are good to go.

Nikhil Sharma 19:28, 20 April 2009 (PST)

Today we started our Growth Assay. I am working with Doug, Patrick, John, and Hank. We followed the basic protocol that we wrote up earlier last week. We took 2 tubes of cells, 280 uL each, and added 10ul water and 60ul KCM to each. We used constructs 11-26, along with DH10B and pBca9495CA-Bca1144 as controls. We plated the samples and placed the plates in the incubator.

Nikhil Sharma 14:31, 22 April 2009 (PST)

Today we pipetted our constructs into a 384 well plate and placed the plate in the Tecan. Doug came in yesterday to pick colonies, and we used these samples to prepare for the assay. We needed two conditions: with arabinose and without. We had to diluted the arabinose by a factor of 1000. We made two solutions of LB with 850 uL each: one with 0.85 uL arabinose and one without. We added 50 uL of each to 16 different wells (total of 32) and then added 1 uL of the vector sample. We took this plate and placed it in the Tecan to be analyzed.

Nikhil Sharma 13:45, 27 April 2009 (PST)

We prepped today to run the full growth assay (5 replicates for each construct). We had to retransform samples 21-24 and streak the other 11, along with our 2 controls. Madhvi helped us out with the controls.

Nikhil Sharma 9:28, 30 April 2009 (PST)

I didn't come into lab yesterday as I had a bad biking accident. My group should have prepared and ran the growth assay.

Nikhil Sharma 15:01, 4 May 2009 (PST)

Apparently, our data from the run my partners last time was bad, so we are re-running the experiment. By the time our data was analyzed, it was too late to start working, so we'll transform on Wednesday.

Nikhil Sharma 16:29, 6 May 2009 (PST)

We transformed all the constructs (11-26) along with the 2 controls today.